In a right angled triangle,
the shortest side = x
the hypotenuse = 2x+1
the remaining side = 2x - 1
Find the length of the hypotenuse

Answer by oscargut(2103)

x^2+(2x-1)^2=(2x+1)^2
x^2+4x^2-4x+1=4x^2+4x+1
x^2-8x=0
x(x-8)=0
then x=8 or x=0
then x=8 and 2x+1=17
Answer the length of the hypotenuse is 17